OpenStudy (anonymous):

lim x→0− 9/x − 9/|x| help would be appreciated!

OpenStudy (anonymous):

I'm getting zero on this again...

OpenStudy (anonymous):

Does 0- mean "approach only from the left"?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

From the negative only side of zero, absolute value of x is always = negative x, because x is neg, so |x| = -x So 9/x - 9/(-x) = 9/x + 9/x = 18/x

OpenStudy (anonymous):

so the limit as you approach 0 of 18/x is infinity (I think)

OpenStudy (anonymous):

Oh.. I got 18 too but I wasnt sure if that was correct but thank you!!

OpenStudy (anonymous):

I don't think the limit is 18 though

OpenStudy (anonymous):

Oh hmm...

OpenStudy (anonymous):

so it would in positive infinity?

OpenStudy (anonymous):

f(x) = 18/x at zero, it's undefined, but close to zero on the negative side, it gets giant quick... yes, positive infinity

OpenStudy (anonymous):

RIght. ah thank you so much!!

OpenStudy (anonymous):

Glad to help :)